OpenStudy (silverlark2):

Without using the calculator, evaluate 3^1/q, given that q = log base 4 of 3.

OpenStudy (anonymous):

\[q=\log_{4} 3\]so,\[3^{\frac{1}{q}}=3^{-q}=3^{-\log_{4}3} \]

OpenStudy (raden):

hint : if q = log base 4 of 3 then 1/q = log base 3 of 4.

OpenStudy (raden):

so, what's would be 3^1/q = 3^( log base 3 of 4) = ... use the property of log :